OpenStudy (anonymous):

inverse functions? what is the inverse function of 2x-5/3

OpenStudy (anonymous):

sorry its f(X)= 2x-5/3

OpenStudy (mathstudent55):

Do you mean \( f(x) = \dfrac{2x - 5}{3} \) ?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

This is how you find the inverse of a function: 1. Replace f(x) with y. 2. Switch x and y. 3. Solve for y. 4. Replace y with \(f^{-1} (x) \).

OpenStudy (mathstudent55):

I'll show you an example with a different, but similar, function. Example: Find the inverse function of \(f(x) = \dfrac{4x - 3}{2} \).

OpenStudy (mathstudent55):

Solution: 1. \(y = \dfrac{4x - 3}{2} \) 2. \(x = \dfrac{4y - 3}{2} \) 3. \(x = \dfrac{4y - 3}{2} \) Multiply both sides by 2: \(2x = 4y - 3 \) Add 3 to both sides: \(2x + 3 = 4y\) Divide both sides by 4: \(\dfrac{2x + 3}{4} = y\) \(y = \dfrac{2x + 3}{4}\) 4. \(f^{-1}(x) = \dfrac{2x + 3}{4}\)

OpenStudy (anonymous):

f\[f ^{-1}(x)=\frac{ 3x+5 }{ 2 }\]

OpenStudy (anonymous):

@mathstudent55 is that right?^

OpenStudy (mathstudent55):

Correct.